OpenStudy (anonymous):

C=4d^-1/3b, Solve for d

OpenStudy (anonymous):

Try to isolate d

OpenStudy (anonymous):

multiply 3b to both sides

OpenStudy (anonymous):

So you'll have 3bc=4d^-1 You'd have to put d as a denominator under 4 (4 being the numerator) Like this:3bc=4/d Then you can multiply 4 to both sides 12bc=d

OpenStudy (jdoe0001):

\(\bf a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\qquad thus \\ \quad \\ \quad \\ c=\cfrac{4d^{-1}}{3b}\implies c=\cfrac{4}{3b}\cdot d^{-1}\implies c=\cfrac{4}{3b}\cdot \cfrac{1}{d^1} \\ \quad \\ c=\cfrac{4}{3bd}\) solve for "d"